Can you please help me solve this equation step by step for g:
.1200=.16(1+g)/30+g
Thank you.
DRS

Answer by stanbon(48568)

0.1200=0.16(1+g)/30+g
Multiply thru by 100 to get:
12 = 16(1+g)/30 + 100g
12 = (8/15)(1+g) + 100g
12-(8/15) = (8/15)g + 100g
(172/15) = (8/15)g + (1500/15)g
Multiply thru by 15 to get:
172 =8g + 1500g
172 = 1508g
g=0.114...
Cheers,
Stan H.